# Surface Tension of a Topological Phase

### v2

@@ Abstract
In interacting topological systems, Landau-like order parameters interplay with the band topology of fermions. The physics of domain formation in such systems can get significantly altered due to the presence of topological fermions. In this work we show that coupling a topological fermionic field to a scalar field can drastically modify the nucleation processes of the scalar field. We find that existence of non-trivial fermionic boundary modes on the nucleating droplets of the scalar field leads to substantial quantum corrections to the surface tension. This leads to an increase in the size of the critical nucleus beyond which unrestricted droplet growth happens. To illustrate the phenomena we devise a minimal model of fermions in a Chern insulating system coupled to a classical Ising field in two spatial dimensions. Using a combination of analytic and numerical methods we conclusively demonstrate that topological phases have a characteristic quantum surface tension. Apart from material systems, our work opens up a host of questions regarding the impact of fermionic topological terms on classical phase transitions and associated criticality.
@@
Introduction: The physics of nucleation, where a droplet having an order parameter value defining the global energy minima grows or shrinks within a metastable phase is one of the defining tenets within the study of phases and phase transitions [1, 2, 3]. These ideas govern a wide range of phenomena spanning condensed matter [4, 5, 6, 7], nuclear physics [8, 9, 10, 11, 12, 13, 14] and high-energy field theories [15, 16, 17, 18, 19, 20, 21]. The conditions on growth and decay of any such nuclei, however, are governed by a competition between surface tension and bulk free-energy density [22, 23, 24]. Microscopic models have been devised to make quantitative predictions and study physical processes such as crystal growth, domain formation and role of thermal excitations and impurity effects [25, 26, 27, 28, 29, 6]. With advances in computational methods and new experimental platforms, many of these ideas have been investigated ranging from material to colloidal systems [30, 31, 32, 33]. While dominantly studied in classical systems, the nucleation physics in presence of quantum fluctuations have been little explored [34]. In particular, the role of topological fermions has not been investigated. Insights from these might be crucial to understand domain formation and nucleation processes in interacting topological insulators where Landau-like order parameters interplay with band topology [35, 36].
@@
Figure 1: Topological droplet: (a) Free-energy density (blue curves) for a two-dimensional classical Ising model coupled to topological fermions in magnetic field $h$ and temperature $T$ where $C$ is Chern number and arrow shows sudden quench. (b) A $C=1$ droplet within a metastable region of $C=-1$. (c) A droplet of size $n$ either grows or shrinks depending on a critical size. This critical size for classical Ising system ($n_{c}^{\rm{cl}}$) increases to $n_{c}^{\rm{Q}}$ when coupled to topological fermions due to droplet edge-modes leading to additional surface tension.
@@
In this work we pose and solve the question – can a topological fermionic field, when coupled to a scalar field, modify the physics of nucleation? We answer this in affirmative, uncovering the role of topological edge-modes in modifying the surface tension and nucleation physics. Such edge-modes are a hallmark of topological phases of matter where they retain their robustness even when the bulk is gapped [37, 38, 39, 40, 41]. While our investigation here is theoretical, such systems of coupled fermionic degrees of freedom with Landau-like scalar order parameters have been of immense interest given their direct applicability to a range of material platforms including those of anomalous quantum Hall systems, spintronics and physics of correlated topological phenomena [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54].
@@
We limit our study to a scalar Ising field in two dimensions which undergoes the standard order-disorder transitions in the magnetic field ($h$) and temperature ($T$) plane (see Fig. 1(a)) where the second order thermal phase transition happens at $T=T_{c}$ [55]. We couple a two dimensional fermionic Chern insulator to the Ising field in such a way that the order parameter of scalar field acts like effective mass for the Dirac fermions. Thus the free-energy minima for $h>0$ and $h<0$ (for $T<T_{c}$) regime also correspond to topological insulating phases (for fermions) albeit with different Chern numbers ($C=\pm 1$). We perform a gedanken experiment where a quench is done from $h<0$ to $h>0$ regime such that the system is in a metastable phase of $C=-1$ with spins down polarized while the global free-energy minima is for $C=+1$ phase with spins up polarized. We now investigate the stability of a bubble of $C=+1$ region created within this metastable phase (see Fig. 1(b)). In absence of any fermionic field it is known that a bubble only beyond a critical size $n_{c}^{\rm{cl}}$ grows, while droplets of size $n<n_{c}^{\rm{cl}}$ shrinks with time. Here, we demonstrate that presence of a topological fermionic field when coupled to a scalar field leads to quantum corrections to surface tension essentially due to topological edge-modes on droplet boundary (see Fig. 1(b)). This quantum surface tension leads to an enhancement of the critical nuclei size to $n_{c}^{\rm{Q}}$ such that $n_{c}^{\rm{Q}}$ is always greater than $n_{c}^{\rm{cl}}$ (see Fig. 1(c)). In fact as we will show such enhancement can be tuned by changing the microscopic parameters of theory, easily achieving $n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}\sim 2$ within experimentally relevant scales.
@@
Model: We consider Bernevig-Hughes-Zhang (BHZ) model [56] of spinless fermions coupled to two-dimensional ferromagnetic classical Ising model on a square lattice, such that the Hamiltonian is
@@
where the spin variable for $i$-th unit cell can assume $s_{i}=\pm 1$ and $J>0$. Each unit cell contains $A$ and $B$ sites having staggered masses ±$\pm\kappa s_{i}$. The hopping strengths between nearest neighboring unit cells along $x$-axis and $y$-axis are ${\large{\eta}}_{ij}=\frac{1}{2}(\sigma_{z}+i\sigma_{x})$ and ${\large{\eta}}_{ij}=\frac{1}{2}(\sigma_{z}+i\sigma_{y})$ respectively where $\sigma_{x}$, $\sigma_{y}$, $\sigma_{z}$ are Pauli matrices and ${\bf{\Psi}}_{i}=\begin{pmatrix}c_{iA}&c_{iB}\end{pmatrix}^{\rm{T}}$ with $c_{iA}$ ($c_{iB}$) being annihilation operator of fermion at site $A$ ($B$) of $i$-th unit cell. $\kappa$ is coupling parameter between the fermions and Ising spins. We keep fermionic filling fixed at half. When $\kappa=1$, an up polarized state with magnetization $m=\frac{1}{L^{2}}\sum_{i}\langle s_{i}\rangle=+1$ corresponds to a fermionic Chern insulator with $C=+1$ and a down polarized state ($m=-1$) leads to $C=-1$. Moreover at $T=0$, $\kappa$ itself can be tuned such that when $20<|\kappa|<$ 2 we have a topological phase with $C=\pm 1$ and a trivial phase with $C=0$ when $|\kappa|>2$ [56, 57, 58]. We will consider periodic boundary conditions throughout our work.
[…]
Quantum correction to surface tension: When $\kappa=0$, the fermions and spins are decoupled and the physics of nucleation under the quench protocol is determined by just the classical Ising fields. Thus, change of free-energy in the formation of a circular cluster of size $n$ is
@@
where $\sigma_{\rm{cl}}>0$ is the surface tension which depends on $J$ and temperature $T$ [2, 27]. When $\kappa=0$, change of free energy $\Delta F_{\rm{classical}}$ is maximum for critical cluster-size
@@
Thus, to reduce the free energy, the cluster grows with time when $n>n_{c}^{\rm{cl}}$ and it shrinks when $n<n_{c}^{\rm{cl}}$, determined by a competition of surface tension with bulk energy-density and a system with a higher surface tension $\sigma_{\rm{cl}}$ would pertain to a larger $n_{c}^{\rm{cl}}$.
Figure 2: Surface tension and critical cluster-size: (a) Change of ground-state energy $\Delta E$ for fermionic system with $\sqrt{n}$ where $n$ is cluster-size. (Inset) Correction to surface tension $\Delta\sigma$ obtained from slope of linear fit (see Eq. (5)) in (a) as a function of $\kappa$, where dotted lines denote quantum critical points $\kappa=-2,0,2$. (b) $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}$ (see Eq. (8)) with $\kappa$ and $J$ at $T=0$, where $n_{c}^{\rm{cl}}$ is critical cluster-size for classical situation ($\kappa=0$). Here, number of unit cells is $L^{2}=4096$. In topological situation ($20<|\kappa|<$ 2), $n_{c}^{\rm{Q}}>n_{c}^{\rm{cl}}$ and relative enhancement $\gamma$ increases with decreasing $J$.
Figure 3: Critical cluster-size and droplet growth: (a) $\nu$ as a function of cluster-size $n$ for classical ($\kappa=0$), topological ($\kappa=1.2$) and trivial ($\kappa=2.5$) situations (see Eq. (9)). $\nu=0$ denotes the critical sizes which are $n_{c}^{\rm{cl}}=145$, $n_{c}^{\rm{Q}}=304$, $n_{c}^{\rm{Q}}=158$ respectively. While in topological situation $n_{c}^{\rm{Q}}>n_{c}^{\rm{cl}}$, in trivial situation $n_{c}^{\rm{Q}}\approx n_{c}^{\rm{cl}}$. (b) Initial and final average local magnetization $\langle s_{x,y}\rangle$ when clusters of sizes, $n=441$ ($n>n_{c}^{\rm{Q}}$), $n=225$ ($n_{c}^{\rm{Q}}>n>n_{c}^{\rm{cl}}$), $n=81$ ($n_{c}^{\rm{cl}}>n$) are evolved. The final time $t=100$ is in units of MC steps/site. Upper (lower) panel shows $\kappa=0$ ($\kappa=1.2$) and $J=0.2$, $T=1/\beta=0.33<T_{c}\approx 0.45$, $h=0.02$, $L^{2}=4096$. Averaging is performed over $10^{3}$ simulations. For classical ($\kappa=0$) situation, the cluster grows when $n>n_{c}^{\rm{cl}}$ but in topological situation ($\kappa=1.2$), growth occurs only when $n>n_{c}^{\rm{Q}}$.
@@
where $\Delta E$ is the quantum contribution from fermionic field which can again have both surface and bulk contributions. The bulk contribution from the fermionic field is zero since in the nucleating droplet there is a flip of the effective mass term $M_{\text{eff}}\equiv\kappa m$ under which the ground state energy of fermionic system at half-filling does not change. The surface contributions to $\Delta E$ can however be non-zero depending on the character of the fermionic phases as we discuss below.
To estimate $\Delta E$ we compare the ground state energies ($h=0,T=0$) between configurations (i) where a square region of size $L^{2}$ has all $s_{i}=-1$ and another (ii) where an internal region of size $l^{2}$ (where $l=\sqrt{n}$) has $s_{i}=+1$ while the rest of $s_{i}=-1$. Variation of $\Delta E$ shows a characteristic $\sqrt{n}$ dependence for a finite $\kappa$ (see Fig. 2(a)), thus leading to a quantum contribution to surface tension ($\Delta\sigma$[59, 60] which can be estimated from the slope as
@@
As shown in inset of Fig. 2(a), $\Delta\sigma\rightarrow 0$ when $\kappa\rightarrow 0$ reflecting the decoupled limit. However $\Delta\sigma$ is significantly larger in topological situation ($20<|\kappa|<$ 2) compared to the trivial situation ($|\kappa|>2$). In the topological situation the droplet necessarily hosts robust edge-modes at the boundary due to a change in the Chern number between the cluster and its background. This imposes a reorganization of the energy-spectrum, causing fermions to move up from the bulk energy levels to mid-gap edge modes. Given a bulk gap of $\sim\Delta_{\rm{bg}}$, the edge state contribution to surface tension can be estimated to be
$\Delta\sigma_{\text{edge}}\sim\frac{\Delta_{\rm{bg}}^{2}}{\hbar v_{F}}$ (6)
where $v_{F}$ is the Fermi velocity of the edge modes. This should be contrasted with the trivial situation where no such edge-mode appears resulting in $\Delta\sigma\rightarrow 0$. Thus $\Delta\sigma$ in topological situation is primarily determined by edge-mode energy per unit length ($\Delta{\mathcal{E}}_{\text{edge}}$) of the boundary of droplet. In general, a nucleating domain of one fermionic phase of matter into another, even both being trivially insulating, has a non-zero surface tension contribution due to perturbative reorganization of all bulk states resulting from domain boundary. Such contributions are, however, non-universal and much smaller than those from topological edge modes. A demarcation of such contributions is discussed in supplemental material (SM) [61].
@@
Interestingly, $\gamma$ behaves non-monotonically with $\kappa$, mirroring the behavior of $\Delta\sigma$. The interplay of Ising spin-exchange scale $J$ and edge-mode contribution to $\Delta\sigma$ allows for further tunability of $\gamma$. At $T=0$, surface tension in classical Ising model is $\sigma_{\rm{cl}}=2J$, implying an increase in $\gamma$ with decreasing $J$ in topological situation ($20<|\kappa|<$ 2). For $J\sim 0.2$, $\kappa\sim 1$, one can tune $\gamma\sim 1.5$ which is upto $50\%$ rise in the size of critical nuclei; while in trivial situation ($|\kappa|>2$), $\gamma\sim 1$, implying no significant change in critical cluster-size (see Fig. 2(b)).
@@
While our discussion of topological surface tension has been done in context of model introduced in Eq. (1), the phenomenology holds true in general. For instance, the Hamiltonian here contains, in addition to bare particle-hole symmetry for fermions, a symmetry where the fermionic spectrum remains invariant under spin flip $s_{i}\rightarrow-s_{i}$. This results into the fermionic volume contribution to free energy $\Delta\epsilon=0$ where $\Delta E\propto\Delta\epsilon n$ (see near Eq. (4)). The quench physics then is solely determined by surface term. However, in absence of this symmetry, when $\Delta\epsilon\neq 0$, topological surface tension still dominantly governs the nucleation physics in topological regime albeit with renormalized $n^{\rm{Q}}_{c}$ and $\gamma$ (see SM [61]). To further study if this physics remains stable under thermal fluctuations, we now analyze the system at finite temperatures.

### v3

# Interfacial Line Energy of a Topological Phase

@@ Abstract
In interacting topological systems, Landau-like order parameters interplay with the band topology of fermions. The physics of domain formation in such systems can get significantly altered due to the presence of topological fermions. In this work we show that coupling a topological fermionic field to a scalar field can drastically modify the nucleation processes. We find that existence of non-trivial fermionic boundary modes on the nucleating droplets of the scalar field leads to substantial quantum corrections to the interface energy. This leads to an increase in the size of the critical nucleus beyond which unrestricted droplet growth happens. To illustrate the phenomena we devise a minimal model of fermions in a Chern insulating system coupled to a classical Ising field in two spatial dimensions. Using a combination of analytic and numerical methods we conclusively demonstrate that topological phases have a characteristic quantum correction to the interfacial line energy. Apart from material systems, our work opens up a host of questions regarding the impact of fermionic topological terms on classical phase transitions and associated criticality.
I Introduction
@@
The physics of nucleation, where a droplet having an order parameter value defining the global energy minima grows or shrinks within a metastable phase is one of the defining tenets within the study of phases and phase transitions [13, 19, 42]. These ideas govern a wide range of phenomena spanning condensed matter [29, 41, 40, 32], nuclear physics [21, 59, 60, 5, 37, 7, 30] and high-energy field theories [26, 27, 25, 64, 43, 38, 39]. The conditions on growth and decay of any such nuclei, however, are governed by a competition between surface tension (interfacial line energy for a one-dimensional interface in a two-dimensional system) and bulk free-energy density [10, 34, 22]. Microscopic models have been devised to make quantitative predictions and study physical processes such as crystal growth, adsorption, domain formation and role of thermal excitations, anisotropy in the domains and impurity effects [9, 4, 49, 1, 2, 3, 55, 50, 40]. With advances in computational methods and new experimental platforms, many of these ideas have been investigated ranging from material to colloidal systems [65, 52, 51, 36, 24, 61]. While dominantly studied in classical systems, the nucleation physics in presence of quantum fluctuations and topological fermions have been little explored [28, 53]. Insights from these might be crucial to understand domain formation and nucleation processes in interacting topological insulators where Landau-like order parameters interplay with band topology [48, 33].
@@
Figure 1: Topological droplet: (a) Free-energy density (blue curves) for a two-dimensional classical Ising model coupled to topological fermions in magnetic field $h$ and temperature $T$ where $C$ is Chern number and arrow shows sudden quench. (b) $C=1$ droplet within a metastable region of C=1C=-1. (c) A droplet of size $n$ either grows or shrinks depending on a critical size. This critical size for classical Ising system ($n_{c}^{\rm{cl}}$) increases to $n_{c}^{\rm{Q}}$ when coupled to topological fermions due to droplet edge-modes leading to additional interfacial line energy.
@@
In this work we pose and solve the question – can a topological fermionic field, when coupled to a scalar field, modify the physics of nucleation? We answer this in affirmative, uncovering the role of topological edge-modes in modifying the interface energy and nucleation physics. Such edge-modes are a hallmark of topological phases of matter where they retain their robustness even when the bulk is gapped [35, 47, 20, 15]. While our investigation here is theoretical, such systems of coupled fermionic degrees of freedom with Landau-like scalar order parameters have been of immense interest given their direct applicability to a range of material platforms including those of anomalous quantum Hall systems, spintronics and physics of correlated topological phenomena [44, 66, 56, 57, 58, 23, 14, 68, 70, 31, 69, 67, 6].
@@
We limit our study to a scalar Ising field in two dimensions which undergoes the standard order-disorder transitions in the magnetic field ($h$) and temperature ($T$) plane (see Fig. 1(a)) where the second order thermal phase transition happens at $T=T_{c}$ [46]. We couple a two dimensional fermionic Chern insulator to the Ising field in such a way that the order parameter of scalar field acts like effective mass for the Dirac fermions. Thus the free-energy minima for $h>0$ and $h<0$ (for $T<T_{c}$) regime also correspond to topological insulating phases (for fermions) albeit with different Chern numbers ($C=\pm 1$). We perform a gedanken experiment where a quench is done from $h<0$ to $h>0$ regime such that the system is in a metastable phase of C=1C=-1 with spins down polarized while the global free-energy minima is for $C=+1$ phase with spins up polarized. We now investigate the stability of a bubble of $C=+1$ region created within this metastable phase (see Fig. 1(b)). In absence of any fermionic field it is known that a bubble only beyond a critical size $n_{c}^{\rm{cl}}$ grows, while droplets of size $n<n_{c}^{\rm{cl}}$ shrinks with time. Here, we demonstrate that presence of a topological fermionic field when coupled to a scalar field leads to quantum corrections to the interface energy per unit length (also known as interfacial line energy) essentially due to topological edge-modes on droplet boundary (see Fig. 1(b)). This quantum correction to the interfacial line energy leads to an enhancement of the critical nuclei size to $n_{c}^{\rm{Q}}$ such that $n_{c}^{\rm{Q}}$ is always greater than $n_{c}^{\rm{cl}}$ (see Fig. 1(c)). In fact as we will show such enhancement can be tuned by changing the microscopic parameters of theory, easily achieving $n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}\sim 2$ within experimentally relevant scales.
II Model
@@
We consider Bernevig-Hughes-Zhang (BHZ) model [11] of spinless fermions coupled to two-dimensional ferromagnetic classical Ising model on a square lattice, such that the Hamiltonian is
@@
where the spin variable for $i$-th unit cell can assume $s_{i}=\pm 1$ and $J>0$. Each unit cell contains $A$ and $B$ sites having staggered masses $\pm\kappa s_{i}$. The hopping strengths between nearest neighboring unit cells along $x$-axis and $y$-axis are ${\large{\eta}}_{ij}=\frac{1}{2}(\sigma_{z}+i\sigma_{x})$ and ${\large{\eta}}_{ij}=\frac{1}{2}(\sigma_{z}+i\sigma_{y})$ respectively where $\sigma_{x}$, $\sigma_{y}$, $\sigma_{z}$ are Pauli matrices and ${\bf{\Psi}}_{i}=\begin{pmatrix}c_{iA}&c_{iB}\end{pmatrix}^{\rm{T}}$ with $c_{iA}$ ($c_{iB}$) being annihilation operator of fermion at site AA ($B$) of $i$-th unit cell. $\kappa$ is coupling parameter between the fermions and Ising spins.
Figure 2: Interface energy and critical cluster-size: (a) Change of ground-state energy $\Delta E$ for fermionic system with $\sqrt{n}$ where $n$ is cluster-size. (Inset) Correction to interfacial line energy $\Delta\sigma$ obtained from slope of linear fit (see Eq. (5)) in (a) as a function of $\kappa$, where dotted lines denote quantum critical points $\kappa=-2,0,2$. (b) $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}$ (see Eq. (8)) with $\kappa$ and $J$ at $T=0$, where $n_{c}^{\rm{cl}}$ is critical cluster-size for classical situation ($\kappa=0$). Here, number of unit cells is $L^{2}=4096$. In topological situation ($0<|\kappa|<2$), $n_{c}^{\rm{Q}}>n_{c}^{\rm{cl}}$ and relative enhancement $\gamma$ increases with decreasing $J$.
We note that for a given $\{s_{i}\}$, the fermionic Hamiltonian in Eq. (1) does not have time-reversal and sub-lattice symmetries. However, it preserves the following symmetries:
  1. 1.
    Spin-fermion flip symmetry: This corresponds to the symmetry operation $s_{i}\rightarrow-s_{i},{\bf{\Psi}}_{i}\rightarrow(-1)^{i_{x}+i_{y}}\sigma_{x}{\bf{\Psi}}_{i}\otimes{\mathcal{K}}$ where $\mathcal{K}$} is the complex conjugation and $(i_{x},i_{y})$ is the position of $i$-th unit cell.
  2. 2.
When $\kappa=1$, an up polarized state with magnetization $m=\frac{1}{L^{2}}\sum_{i}\langle s_{i}\rangle=+1$ corresponds to a fermionic Chern insulator with $C=+1$ and a down polarized state (m=1m=-1) leads to C=1C=-1. Moreover at $T=0$, $\kappa$ itself can be tuned such that when $0<|\kappa|<2$ we have a topological phase with $C=\pm 1$ and a trivial phase with $C=0$ when $|\kappa|>2$ [11, 12, 8]. We will consider periodic boundary conditions throughout our work.
Let us now consider the system being subjected to a sudden quench of external field from $h<0$ to $h>0$ (keeping $|h|$ fixed) at a fixed temperature $T<T_{c}$. The system thus forms a metastable state of m=1m=-1, whose stability to small droplets of $m=+1$ (the global minima) can be analyzed, as we discuss below.
III Topological interface energy
When $\kappa=0$, the fermions and spins are decoupled and the physics of nucleation under the quench protocol is determined by just the classical Ising fields. Thus, change of free-energy in the formation of a circular cluster of size $n$ is
@@ III Topological interface energy
where $\sigma_{\rm{cl}}>0$ is the interfacial line energy which depends on $J$ and temperature $T$ [9, 4, 19, 17]. When $\kappa=0$, change of free energy $\Delta F_{\rm{classical}}$ is maximum for critical cluster-size
@@
Thus, to reduce the free energy, the cluster grows with time when $n>n_{c}^{\rm{cl}}$ and it shrinks when $n<n_{c}^{\rm{cl}}$, determined by a competition of the interfacial line energy with the bulk energy-density and a system with a higher $\sigma_{\rm{cl}}$ would pertain to a larger $n_{c}^{\rm{cl}}$.
@@
where $\Delta E$ is the quantum contribution from fermionic field which can again have both the interface and bulk contributions. However, due to the spin-fermion flip symmetry, the energy eigenvalues $E_{j}$ of bulk spectrum for a given $\{s_{i}\}$ are independent of the sign of $\kappa$, i.e. $E_{j}(\kappa)=E_{j}(-\kappa)$. This also reflects into a spin-flip symmetry as $E_{j}(\kappa,\{s_{i}\})=E_{j}(\kappa,\{-s_{i}\})$. This guarantees that under $s_{i}\to-s_{i}$, the change in the bulk energy density of fermionic system $\Delta\epsilon=0$. In other words, the bulk contribution from the fermionic field is zero since in the nucleating droplet there is a flip of the effective mass term $M_{\text{eff}}\equiv\kappa m$ under which the ground state energy of fermionic system at half-filling does not change. The interfacial contribution to $\Delta E$ can however be non-zero depending on the character of the fermionic phases as we discuss below.
To estimate $\Delta E$ we compare the ground state (i.e. at half-filling) energies ($h=0,T=0$) between configurations (i) where a square region of size $L^{2}$ has all $s_{i}=-1$ and another (ii) where an internal square-shaped region of size $l^{2}$ (where there are $l=\sqrt{n}$ unit cells along both $x$ and $y$-axis) has $s_{i}=+1$ while the rest of $s_{i}=-1$. Variation of $\Delta E$ shows a characteristic $\sqrt{n}$ dependence for a finite $\kappa$ (see Fig. 2(a)), thus leading to a quantum contribution to the interfacial line energy ($\Delta\sigma$[63, 62] which can be estimated from the slope as
@@
As shown in inset of Fig. 2(a), $\Delta\sigma\rightarrow 0$ when $\kappa\rightarrow 0$ reflecting the decoupled limit. However $\Delta\sigma$ is significantly larger in topological situation ($0<|\kappa|<2$) compared to the trivial situation ($|\kappa|>2$). In the topological situation the droplet necessarily hosts robust edge-modes at the boundary due to a change in the Chern number between the cluster and its background. This imposes a reorganization of the energy-spectrum, causing fermions to move up from the bulk energy levels to mid-gap edge modes. Given a bulk gap of $\sim\Delta_{\rm{bg}}$, the correction to the interfacial line energy in the topological situation is primarily determined by the edge-mode energy per unit length of the boundary of the droplet (see Appendix A):
$\Delta{\mathcal{E}}_{\rm{edge}}\sim\frac{\Delta_{\rm{bg}}^{2}}{16\pi\hbar v_{F}}$ (6)
where $v_{F}$ is the Fermi velocity of the edge modes. This should be contrasted with the trivial situation where no such edge-mode appears resulting in $\Delta\sigma\rightarrow 0$. Thus, the role of edge-modes as the primary source of the interfacial line energy in the topological phase is established. In general, a nucleating domain of one fermionic phase of matter into another, even both being trivially insulating, has a non-zero interface energy contribution due to perturbative reorganization of all bulk states resulting from domain boundary. Such contributions are, however, non-universal and much smaller than those from topological edge modes. A demarcation of such contributions is discussed in Appendix A.
Figure 3: Local magnetization in coupled spin-fermion Monte Carlo simulation: Local magnetization $s_{x,y}$ for unit cell with position $(x,y)$ at times $t=0,20,40$ (in units of Monte Carlo steps per site) in a typical realization of the coupled spin-fermion exact Monte Carlo simulation (see Eq. (10) and Eq. (11)) for $\kappa=1.2$. Here, the initial cluster-size with up-spins is $n=441$. The parameters chosen are: $J=0.2$, $T=1/\beta=0.33$ (where $T_{c}\approx 0.45$), $h=0.02$ and number of unit cells in square lattice $L^{2}=4096$. We observe the growth of the cluster.
Figure 4: Local magnetization in a typical realization: Local magnetization $s_{x,y}$ for unit cell with position $(x,y)$ at times $t=0,20,100$ (in units of Monte Carlo steps per site) in a typical realization for $\kappa=0$ and $\kappa=1.2$ where initial cluster-sizes with up-spins are $n=441$, $n=225$ and $n=81$. The parameters chosen are: $J=0.2$, $T=1/\beta=0.33$ (where $T_{c}\approx 0.45$), $h=0.02$. Number of unit cells in square lattice is $L^{2}=4096$. The cluster with initial size $n=441$ grows in both the decoupled ($\kappa=0$) and topological ($\kappa=1.2$) situations, while the cluster with $n=81$ shrinks in both the situations. When $n=225$, we observe the growth of the cluster for $\kappa=0$ and the shrinkage of the cluster for $\kappa=1.2$.
@@
Interestingly, $\gamma$ behaves non-monotonically with $\kappa$, mirroring the behavior of $\Delta\sigma$. The interplay of Ising spin-exchange scale $J$ and edge-mode contribution to $\Delta\sigma$ allows for further tunability of $\gamma$. At $T=0$, the interfacial line energy in classical Ising model with square-shaped droplet boundary is $\sigma_{\rm{cl}}\sim 2J$, while the exact value of $\sigma_{\rm{cl}}$ depends on the anisotropy of the droplet and the direction of its growth [9, 4]. We observe an increase in $\gamma$ with decreasing $J$ in topological situation ($0<|\kappa|<2$). For $J\sim 0.2$, $\kappa\sim 1$, one can tune $\gamma\sim 1.5$ which is upto $50\%$ rise in the size of critical nuclei; while in trivial situation ($|\kappa|>2$), $\gamma\sim 1$, implying no significant change in critical cluster-size (see Fig. 2(b)).
@@
While our discussion of topological interface energy has been done in context of model introduced in Eq. (1), the phenomenology holds true in general. Due to the spin-fermion flip symmetry for the model in Eq. (1), the fermionic volume contribution to free energy $\Delta\epsilon=0$ where $\Delta E\propto\Delta\epsilon n$ (see near Eq. (4)). The quench physics then is solely determined by the interfacial contribution. However, in absence of this symmetry, when $\Delta\epsilon\neq 0$, the topological interface energy contribution still dominantly governs the nucleation physics in topological regime albeit with renormalized $n^{\rm{Q}}_{c}$ and $\gamma$ (see Appendix B). To further study if this physics remains stable under thermal fluctuations, we now analyze the system at finite temperatures.
IV Growth of clusters at finite temperature